4. Find the square roots of the following numbers by the Prime Factorisation Method.   vii. 5929  

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student


5929 = 7×7×11×11
⇒ 5929 =(7×7)×(11×11)
⇒ 5929 =(7×11)×(7×11)
⇒ 5929 = (7×11)2
⇒√5929=7×11=77